# Query-Efficient Zeroth-Order Algorithms for Nonconvex Optimization

### v1

@@ 7  Conclusion
In this research, we study a general optimization problem with black-box constraints. We reformulate it as a min-max problem, and then apply zeroth-order optimization (ZO) methods to solve it using the input-output information. Specifically, by integrating block updates with gradient descent ascent (GDA), we develop two novel algorithms, called ZOB-GDA and ZOB-SGDA, which achieve efficiency in both single-step gradient estimation and the overall query complexity. Our theoretical results demonstrate that ZOB-GDA achieves the same query complexity as its first-order counterpart with an additional dimension-dependent factor, and ZOB-SGDA enjoys the best-known complexity bound. In addition, our numerical experiments validate the superior performance of our algorithms. However, our work on block updates in constrained ZO is just a beginning. There are still several open challenges. First, while the block update framework is a broadly applicable technique for improving single-step efficiency, its integration with other primal–dual algorithms requires more study. Second, although we anticipate that the benefits of block updates in stochastic constrained ZO will be more limited than in deterministic settings, rigorous validation requires further investigation.
@@
Define the filtration: $\mathcal{F}_{k}=\sigma(x_{0},y_{0},\mathcal{I}_{0},\cdots,i_{k-1},x_{k},y_{k})$. Then, we provide the following lemma to bound the one-step drift of $\Phi_{1/2L}(x_{k})$.
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\Delta_{k}\leq$ $\alpha\Lambda_{0}^{2}(B+2)+\frac{R_{y}^{2}}{2\beta B}+\frac{\Delta_{0}}{K}.$
@@
$\min_{k\leq K-1}\mathbb{E}\left[\left\|\Phi_{1/2L}(x_{k})\right\|\right]\leq\left(\mathcal{O}\left(\frac{N}{\alpha K}\right)+\epsilon_{c}^{2}\right)^{1/2}\leq\mathcal{O}\left(\sqrt{\frac{N}{\alpha K}}\right)+\epsilon_{c}.$
@@
which is derived from Davis and Drusvyatskiy, (2019), to get
@@
$\sum_{k=jB}^{(j+1)B-1}\Delta_{k}\leq\;$ $\alpha\Lambda_{0}^{2}(B^{2}+B)+\frac{R_{y}^{2}}{2\beta}+\mathbb{E}\left[f(x_{jB+B},y_{jB+B})-f(x_{jB},y_{jB})\right].$
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\Delta_{k}\leq\;$ $\alpha\Lambda_{0}^{2}(B+1)+\frac{R_{y}^{2}}{2\beta B}+\frac{1}{K}\mathbb{E}\left[f(x_{K},y_{K})-f(x_{0},y_{0})\right].$
@@
$d(y,z)=\min_{x\in\mathbb{R}^{d_{x}}}K(x,y;z),\;\;m(z)=\min_{x\in\mathbb{R}^{d_{x}}}\max_{y\in\mathcal{Y}}K(x,y;z),$
$h(x,z)=\max_{y\in\mathcal{Y}}K(x,y;z),\;\;x(y,z)=\arg\min_{x\in\mathbb{R}^{d_{x}}}K(x,y;z),$
$x^{*}(z)=\arg\min_{x\in\mathbb{R}^{d_{x}}}h(x,z),\;\;\mathcal{Y}(z)=\arg\max_{y\in\mathcal{Y}}d(y,z),$
$y_{+}(z_{k})=\mathcal{P}_{\mathcal{Y}}[y_{k}+\beta\nabla_{y}K(x(y_{k},z_{k}),y_{k},z_{k})].$
Note that $\mathcal{Y}(z)$ is a set, and we use $y(z)$ to denote an arbitrary element in $\mathcal{Y}(z)$. Recall that we assume $f(x,y)$ is $L$-smooth in $x$ and $y$. Then, if $p>L$, $K(x,y;z)$ is (pL)(p-L)-strongly convex in $x$ and smooth in $x$ with a constant $(L+p)$. We define the potential function:
$\phi(x,y,z)=K(x,y;z)-2d(y,z)+2m(z).$
@@
Lemma B.2 follows from the results in (Zhang et al.,, 2020, Lemma B.2). Therefore, we omit its proof here.
@@
Based on Danskin’s Theorem, we have $\nabla_{y}d(y,z)=\nabla_{y}K(x(y,z),y;z)=\nabla_{y}f(x(y,z),y)$. Then, we have for any $y,y^{\prime}\in\mathcal{Y}$
@@
$=\;$ $\|\nabla_{y}K(x(y,z),y;z)-\nabla_{y}K(x(y^{\prime},z),y^{\prime};z)\|$
$\leq\;$ $\|\nabla_{y}K(x(y,z),y;z)-\nabla_{y}K(x(y,z),y^{\prime};z)\|$
$+\|\nabla_{y}K(x(y,z),y^{\prime};z)-\nabla_{y}K(x(y^{\prime},z),y^{\prime};z)\|$
@@
where the third step follows from the $L$-smoothness of $K(x,y;z)$ in $y$, and the last step follows from Lemma B.2. ∎
@@
Denote $\mathcal{F}_{k}=\sigma(x_{0},y_{0},\mathcal{I}_{0},\cdots,i_{k-1},x_{k},y_{k})$ as a filtration. By Lemma 3.10 in Zhang and Luo, (2020), we have
$\|x_{k}-x(y_{k},z_{k})\|\leq$ $\frac{1}{\alpha(p-L)}\left\|x_{k}-\mathcal{P}_{\mathcal{X}}[x_{k}-\alpha\nabla_{x}K(x_{k},y_{k};z_{k})]\right\|$ (20)
$=$ $\frac{1}{\alpha(p-L)}\left\|\alpha\nabla_{x}K(x_{k},y_{k};z_{k})\right\|,$
@@
$\leq\;$ $2\|x_{k+1}-x_{k}\|^{2}+\frac{2}{\alpha^{2}(p-L)^{2}}\left\|\alpha\nabla_{x}K(x_{k},y_{k};z_{k})\right\|^{2}$
@@
$=\;$ $\left\|\mathcal{P}_{\mathcal{Y}}[y_{k}-\beta\cdot\nabla_{y}K(x_{k},y_{k};z_{k})]-\mathcal{P}_{\mathcal{Y}}[y_{k}-\beta\cdot\nabla_{y}K(x(y_{k},z_{k}),y_{k};z_{k})]\right\|^{2}$
$\leq\;$ $\beta^{2}\|\nabla_{y}K(x(y_{k},z_{k}),y_{k};z_{k})-\nabla_{y}K(x_{k},y_{k};z_{k})\|^{2}$
@@
where in the first step we used the non-expansiveness of projection operations and in the third step we used the $L$-smoothness of $K(x,y;z)$ in $x$. Then, taking the expectation of the above inequality leads to
@@
$\beta(p-L)\|x^{*}(z_{k})-x(y_{+}(z_{k}),z_{k})\|^{2}\leq(1+\beta L+\beta L\sigma_{2})D_{y}\|y_{k}-y_{+}(z_{k})\|.$
Lemma B.6 comes from (Zhang et al.,, 2020, Lemma B.10).
@@
Step 1: Derive a Bound for the Stationarity Measure. Recall that the stationarity measure is defined as $M(x,y)=\min\{\|\Phi(x)\|,\|\mathfrak{g}(x,y)\|\}$. Here we analyze the convergence of $\|\mathfrak{g}(x,y)\|$. Also, recall the definition of $\mathfrak{g}(x,y)$:
@@
$=\;$ $\mathbb{E}\left[\|G_{x}(x_{k},y_{k};z_{k})+\nabla_{x}K(x_{k},y_{k}z_{k})-G_{x}(x_{k},y_{k};z_{k})-p(x_{k}-z_{k})\|^{2}\right]$
$\leq\;$ $\mathbb{E}\left[3\|G_{x}(x_{k},y_{k};z_{k})\|^{2}+3\|G_{x}(x_{k},y_{k};z_{k})-\nabla_{x}K(x_{k},y_{k};z_{k})\|^{2}+3p^{2}\|x_{k}-z_{k}\|^{2}\right]$
@@
$+\langle\nabla_{y}K(x_{k+1},y_{k};z_{k}),\;y_{k}-y_{k+1}\rangle-\frac{L^{2}r_{k}^{2}d_{x}}{8N}.$
@@
By the update of $x$, we use the smoothness of $K$ to get
$K(x_{k+1},y_{k};z_{k})-K(x_{k},y_{k};z_{k})$ (22)
$\leq$ $\langle\nabla_{x}K(x_{k},y_{k};z_{k}),\;x_{k+1}-x_{k}\rangle+\frac{p+L}{2}\|x_{k+1}-x_{k}\|^{2}$
@@
$+\langle\nabla_{x}K(x_{k},y_{k};z_{k})-G_{x}(x_{k},y_{k};z_{k}),\;x_{k+1}-x_{k}\rangle$
@@
where in the second inequality we applied AM-GM inequality and the fact that only the entries of $\mathcal{I}_{k}$ in $x_{k+1}-x_{k}$ are nonzero. Similarly, we can use the $L$-smoothness of $K(x,y;z)$ to get
$K(x_{k+1},y_{k};z_{k})-K(x_{k+1},y_{k+1};z_{k})$
$\geq\;$ $\langle\nabla_{y}K(x_{k+1},y_{k};z_{k}),\;y_{k}-y_{k+1}\rangle-\frac{L}{2}\|y_{k+1}-y_{k}\|^{2}.$ (23)
@@
We further provide the following two lemmas to characterize the one-step drift of $d(y,z)$ and $m(z)$ in ZOB-SGDA. Their proofs follow from (Zhang et al.,, 2020, Lemma B.6 and Lemma B.7).
@@
$\geq\;$ $\langle\nabla_{y}K(x(y_{k},z_{k}),y_{k};z_{k}),\;y_{k+1}-y_{k}\rangle-\frac{L_{d}}{2}\|y_{k+1}-y_{k}\|^{2}$
@@
$=\;$ $\langle\nabla_{y}K(x(y_{k},z_{k}),y_{k};z_{k}),y_{k+1}-y_{k}\rangle-\frac{L_{d}}{2}\|y_{k+1}-y_{k}\|^{2},$
where in the second step we used $\nabla_{y}d(y_{k},z_{k})=\nabla_{y}K(x(y_{k},z_{k}),y_{k};z_{k})$. Also, we have
@@
Using Kakutoni’s Theorem, we have
@@
$+\frac{p}{2\gamma}\|z_{k+1}-z_{k}\|^{2}+\langle\nabla_{y}K(x_{k+1},y_{k};z_{k}),y_{k+1}-y_{k}\rangle$
$+2\langle\nabla_{y}K(x(y_{k},z_{k}),y_{k};z_{k})-\nabla_{y}K(x_{k+1},y_{k};z_{k}),\;y_{k+1}-y_{k}\rangle$
@@
$\langle y_{k+1}-(y_{k}+\beta\nabla_{y}K(x_{k},y_{k};z_{k})),\;y_{k+1}-y\rangle\leq 0,$
@@
$\langle\nabla_{y}K(x_{k},y_{k};z_{k}),y_{k+1}-y_{k}\rangle\geq\frac{1}{\beta}\|y_{k+1}-y_{k}\|^{2}.$
@@
$\langle\nabla_{y}K(x_{k+1},y_{k};z_{k}),y_{k+1}-y_{k}\rangle$
$=\;$ $\langle\nabla_{y}K(x_{k},y_{k};z_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}K(x_{k+1},y_{k};z_{k})-\nabla_{y}K(x_{k},y_{k};z_{k}),y_{k+1}-y_{k}\rangle$
@@
where the first inequality follows from AM-GM inequality and the smoothness of $K$. Then, we can further get
@@
$+2\langle\nabla_{y}K(x(y_{k},z_{k}),y_{k};z_{k})-\nabla_{y}K(x_{k+1},y_{k};z_{k}),\;y_{k+1}-y_{k}\rangle$
@@
$\mathbb{E}\left[2\langle\nabla_{y}K(x(y_{k},z_{k}),y_{k};z_{k})-\nabla_{y}K(x_{k+1},y_{k};z_{k}),\;y_{k+1}-y_{k}\rangle\right]$ (27)
@@
where the first step follows from the smoothness of $K(x,y;z)$ and the third step follows from Lemma B.4. In the last step, we used the fact that $p>L$. Taking the expectation on both sides of (LABEL:eq:potential_bound1) and combining it with the bounds in (LABEL:eq:bound_term1) and (LABEL:eq:bound_term2), we can get
@@
Let $t_{1}=384pD_{y}\frac{1+\beta L+\beta L\sigma_{2}}{p-L}$. Then, using Lemma B.6, we have
@@
$\leq\;$ $384p\gamma\beta D_{y}\frac{1+\beta L+\beta L\sigma_{2}}{\beta(p-L)}\|y_{k}-y_{+}(z_{k})\|,$
@@
$\leq\;$ $384\gamma^{2}D_{y}\frac{1+\beta L+\beta L\sigma_{2}}{\beta(p-L)}\|y_{k}-y_{+}(z_{k})\|$
@@
$\leq\;$ $384p\alpha\gamma D_{y}\frac{1+\beta L+\beta L\sigma_{2}}{\beta(p-L)}\|y_{k}-y_{+}(z_{k})\|$
@@
By $\gamma\leq\frac{1}{\sqrt{KN}}$ and $\sum_{k=0}^{K-1}r_{k}^{2}=\frac{1}{b}$, we can get $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\mathfrak{g}(x_{k},y_{k})\|^{2}\right]\leq\mathcal{O}\left(\frac{N}{K}+\sqrt{\frac{N}{K}}\right)$, which leads to
@@
Substituting $\gamma=\frac{1}{\sqrt{KN}}$ and $\sum_{k=0}^{K-1}r_{k}^{2}\leq\frac{1}{b}$ into the above inequality, we can get
@@
We provide the definition of a critical KKT point below.
Definition C.1.
We say a point $x^{*}$ is a critical KKT point of problem (1) if $c_{j}(x^{*})\leq 0,\forall j\in\mathcal{J}$ and there exists $y^{*}\in\mathcal{Y}$ satisfying $y^{*}\geq 0$ and
$y^{*}(j)c_{j}(x^{*})=0,\forall j\in\mathcal{J},$ (30)
$\nabla_{x}h(x^{*})+(y^{*})^{T}\nabla_{x}c(x^{*})=0.$ (31)
We restate Lemma 5.1 as follows.
Lemma C.1.
Suppose that $(x,y)\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$ is a stationery point of $f(x,y)$ satisfying $\|\mathfrak{g}(x,y)\|=0$ and $y<\overline{y}$. Then, $x$ is a critical KKT point of problem (1).
Proof of Lemma 5.1.
Recall that $\|\mathfrak{g}(x,y)\|=0$ for some $(x,y)\in\mathbb{R}^{d_{x}}\times\mathcal{Y}$ implies that
$\begin{pmatrix}\mathfrak{g}_{x}(x,y)\\ \mathfrak{g}_{y}(x,y)\end{pmatrix}$=\; $\begin{pmatrix}\nabla_{x}f(x,y)\\ (1/\beta)\left(y-\mathcal{P}_{\mathcal{Y}}\left[y+\beta\nabla_{y}f(x,y)\right]\right)\end{pmatrix}$
$=\;$ $\begin{pmatrix}\nabla_{x}h(x)+y^{T}\nabla_{x}c(x)\\ (1/\beta)\left(y-\mathcal{P}_{\mathcal{Y}}\left[y+\beta c(x)\right]\right)\end{pmatrix}$
$=\;$ $0.$
The first condition $\nabla_{x}h(x)+y^{T}\nabla_{x}c(x)=0$ implies (31) directly. The second condition implies that $y-\mathcal{P}_{\mathcal{Y}}\left[y+\beta c(x)\right]=0$. Then, due to that $y<\overline{y}$, one of the following two cases holds: (1) $c(x)=0$; (2) $y=0$ and $c(x)\leq 0$. Both cases can lead to $c(x)\leq 0$ and (30).

### v2

@@ 7  Conclusion
In this research, we study a general optimization problem with black-box constraints. We reformulate it as a min-max problem, and then apply zeroth-order optimization (ZO) methods to solve it using the input-output information. Specifically, by integrating block updates with gradient descent ascent (GDA), we develop two novel algorithms, called ZOB-GDA and ZOB-SGDA, which achieve efficiency in both single-step gradient estimation and the overall query complexity. Our theoretical results demonstrate that ZOB-GDA achieves the same query complexity bound as its first-order counterpart with an additional dimension-dependent factor, and ZOB-SGDA enjoys the best-known complexity bound. In addition, our numerical experiments validate their superior performance. However, our work on block updates in constrained ZO is just a beginning. There are still several open challenges. First, while the block update framework is a broadly applicable technique for improving single-step efficiency, its integration with other primal–dual algorithms requires more study. Second, although we anticipate that the benefits of block updates in stochastic constrained ZO will be more limited than in deterministic settings, rigorous validation requires further investigation.
@@
Define the filtration: $\mathcal{F}_{k}=\sigma(x_{0},y_{0},\mathcal{I}_{0},\cdots,\mathcal{I}_{k-1},x_{k},y_{k})$. Then, we provide the following lemma to bound the one-step drift of $\Phi_{1/2L}(x_{k})$.
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\Delta_{k}\leq$ $\alpha\Lambda_{0}^{2}(B+1)+\frac{R_{y}^{2}}{2\beta B}+\frac{\Delta_{0}}{K}.$
@@
$\min_{k\leq K-1}\mathbb{E}\left[\left\|\nabla\Phi_{1/2L}(x_{k})\right\|\right]\leq\left(\mathcal{O}\left(\frac{N}{\alpha K}\right)+\epsilon_{c}^{2}\right)^{1/2}\leq\mathcal{O}\left(\sqrt{\frac{N}{\alpha K}}\right)+\epsilon_{c}.$
@@
which is derived from Davis and Drusvyatskiy (2019), to get
@@
$\sum_{k=jB}^{(j+1)B-1}\Delta_{k}\leq\;$ $\alpha\Lambda_{0}^{2}B^{2}+\frac{R_{y}^{2}}{2\beta}+\mathbb{E}\left[f(x_{jB+B},y_{jB+B})-f(x_{jB},y_{jB})\right].$
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\Delta_{k}\leq\;$ $\alpha\Lambda_{0}^{2}B+\frac{R_{y}^{2}}{2\beta B}+\frac{1}{K}\mathbb{E}\left[f(x_{K},y_{K})-f(x_{0},y_{0})\right].$
@@
$d(y,z)=\min_{x\in\mathbb{R}^{d_{x}}}H(x,y;z),\;\;m(z)=\min_{x\in\mathbb{R}^{d_{x}}}\max_{y\in\mathcal{Y}}H(x,y;z),$
$\psi(x,z)=\max_{y\in\mathcal{Y}}H(x,y;z),\;\;x(y,z)=\arg\min_{x\in\mathbb{R}^{d_{x}}}H(x,y;z),$
$x^{*}(z)=\arg\min_{x\in\mathbb{R}^{d_{x}}}\psi(x,z),\;\;\mathcal{Y}(z)=\arg\max_{y\in\mathcal{Y}}d(y,z),$
$y_{+}(z_{k})=\mathcal{P}_{\mathcal{Y}}[y_{k}+\beta\nabla_{y}H(x(y_{k},z_{k}),y_{k},z_{k})].$
Note that $\mathcal{Y}(z)$ is a set, and we use $y(z)$ to denote an arbitrary element in $\mathcal{Y}(z)$. Recall that we assume $f(x,y)$ is $L$-smooth in $x$ and $y$. Then, if $p>L$, $H(x,y;z)$ is (pL)(p-L)-strongly convex in $x$ and smooth in $x$ with a constant $(L+p)$. We define the potential function:
$\phi(x,y,z)=H(x,y;z)-2d(y,z)+2m(z).$
@@
Lemma B.2 follows from the results in (Zhang et al., 2020, Lemma B.2). Therefore, we omit its proof here.
@@
Based on Danskin’s Theorem, we have $\nabla_{y}d(y,z)=\nabla_{y}H(x(y,z),y;z)=\nabla_{y}f(x(y,z),y)$. Then, we have for any $y,y^{\prime}\in\mathcal{Y}$
@@
$=\;$ $\|\nabla_{y}H(x(y,z),y;z)-\nabla_{y}H(x(y^{\prime},z),y^{\prime};z)\|$
$\leq\;$ $\|\nabla_{y}H(x(y,z),y;z)-\nabla_{y}H(x(y,z),y^{\prime};z)\|$
$+\|\nabla_{y}H(x(y,z),y^{\prime};z)-\nabla_{y}H(x(y^{\prime},z),y^{\prime};z)\|$
@@
where the third step follows from the $L$-smoothness of $H(x,y;z)$ in $y$, and the last step follows from Lemma B.2. ∎
@@
Denote $\mathcal{F}_{k}=\sigma(x_{0},y_{0},\mathcal{I}_{0},\cdots,i_{k-1},x_{k},y_{k})$ as a filtration. By Lemma 3.10 in Zhang and Luo (2020), we have
$\|x_{k}-x(y_{k},z_{k})\|\leq$ $\frac{1}{\alpha(p-L)}\left\|x_{k}-\mathcal{P}_{\mathcal{X}}[x_{k}-\alpha\nabla_{x}H(x_{k},y_{k};z_{k})]\right\|$ (21)
$=$ $\frac{1}{\alpha(p-L)}\left\|\alpha\nabla_{x}H(x_{k},y_{k};z_{k})\right\|,$
@@
$\leq\;$ $2\|x_{k+1}-x_{k}\|^{2}+\frac{2}{\alpha^{2}(p-L)^{2}}\left\|\alpha\nabla_{x}H(x_{k},y_{k};z_{k})\right\|^{2}$
@@
$=\;$ $\left\|\mathcal{P}_{\mathcal{Y}}[y_{k}+\beta\cdot\nabla_{y}H(x_{k},y_{k};z_{k})]-\mathcal{P}_{\mathcal{Y}}[y_{k}+\beta\cdot\nabla_{y}H(x(y_{k},z_{k}),y_{k};z_{k})]\right\|^{2}$
$\leq\;$ $\beta^{2}\|\nabla_{y}H(x(y_{k},z_{k}),y_{k};z_{k})-\nabla_{y}H(x_{k},y_{k};z_{k})\|^{2}$
@@
where in the first step we used the non-expansiveness of projection operations and in the third step we used the $L$-smoothness of $H(x,y;z)$ in $x$. Then, taking the expectation of the above inequality leads to
@@
$\beta(p-L)\|x^{*}(z_{k})-x(y_{+}(z_{k}),z_{k})\|^{2}\leq(1+\beta L+\beta L\sigma_{2})R_{y}\|y_{k}-y_{+}(z_{k})\|.$
Lemma B.6 comes from (Zhang et al., 2020, Lemma B.10).
@@
Step 1: Derive a Bound for the Stationarity Measure. For ZOB-SGDA, we analyze the convergence of $\|\mathfrak{g}(x,y)\|$. Recall the definition of $\mathfrak{g}(x,y)$:
@@
$=\;$ $\mathbb{E}\left[\|G_{x}(x_{k},y_{k};z_{k})+\nabla_{x}H(x_{k},y_{k};z_{k})-G_{x}(x_{k},y_{k};z_{k})-p(x_{k}-z_{k})\|^{2}\right]$
$\leq\;$ $\mathbb{E}\left[3\|G_{x}(x_{k},y_{k};z_{k})\|^{2}+3\|G_{x}(x_{k},y_{k};z_{k})-\nabla_{x}H(x_{k},y_{k};z_{k})\|^{2}+3p^{2}\|x_{k}-z_{k}\|^{2}\right]$
@@
$+\langle\nabla_{y}H(x_{k+1},y_{k};z_{k}),\;y_{k}-y_{k+1}\rangle-\frac{L^{2}r_{k}^{2}d_{x}}{8N}.$
@@
By the update of $x$, we use the smoothness of $H$ to get
$H(x_{k+1},y_{k};z_{k})-H(x_{k},y_{k};z_{k})$ (23)
$\leq$ $\langle\nabla_{x}H(x_{k},y_{k};z_{k}),\;x_{k+1}-x_{k}\rangle+\frac{p+L}{2}\|x_{k+1}-x_{k}\|^{2}$
@@
$+\langle\nabla_{x}H(x_{k},y_{k};z_{k})-G_{x}(x_{k},y_{k};z_{k}),\;x_{k+1}-x_{k}\rangle$
@@
where in the second inequality we applied AM-GM inequality and the fact that only the entries of $\mathcal{I}_{k}$ in $x_{k+1}-x_{k}$ are nonzero. Similarly, we can use the $L$-smoothness of $H(x,y;z)$ to get
$H(x_{k+1},y_{k};z_{k})-H(x_{k+1},y_{k+1};z_{k})$
$\geq\;$ $\langle\nabla_{y}H(x_{k+1},y_{k};z_{k}),\;y_{k}-y_{k+1}\rangle-\frac{L}{2}\|y_{k+1}-y_{k}\|^{2}.$ (24)
@@
We further provide the following two lemmas to characterize the one-step drift of $d(y,z)$ and $m(z)$ in ZOB-SGDA. Their proofs follow from (Zhang et al., 2020, Lemma B.6 and Lemma B.7).
@@
$\geq\;$ $\langle\nabla_{y}H(x(y_{k},z_{k}),y_{k};z_{k}),\;y_{k+1}-y_{k}\rangle-\frac{L_{d}}{2}\|y_{k+1}-y_{k}\|^{2}$
@@
$=\;$ $\langle\nabla_{y}H(x(y_{k},z_{k}),y_{k};z_{k}),y_{k+1}-y_{k}\rangle-\frac{L_{d}}{2}\|y_{k+1}-y_{k}\|^{2},$
where in the second step we used $\nabla_{y}d(y_{k},z_{k})=\nabla_{y}H(x(y_{k},z_{k}),y_{k};z_{k})$. Also, we have
@@
For any $y(z)\in\mathcal{Y}(z)$, we have
@@
$+\frac{p}{2\gamma}\|z_{k+1}-z_{k}\|^{2}+\langle\nabla_{y}H(x_{k+1},y_{k};z_{k}),y_{k+1}-y_{k}\rangle$
$+2\langle\nabla_{y}H(x(y_{k},z_{k}),y_{k};z_{k})-\nabla_{y}H(x_{k+1},y_{k};z_{k}),\;y_{k+1}-y_{k}\rangle$
@@
$\langle y_{k+1}-(y_{k}+\beta\nabla_{y}H(x_{k},y_{k};z_{k})),\;y_{k+1}-y\rangle\leq 0,$
@@
$\langle\nabla_{y}H(x_{k},y_{k};z_{k}),y_{k+1}-y_{k}\rangle\geq\frac{1}{\beta}\|y_{k+1}-y_{k}\|^{2}.$
@@
$\langle\nabla_{y}H(x_{k+1},y_{k};z_{k}),y_{k+1}-y_{k}\rangle$
$=\;$ $\langle\nabla_{y}H(x_{k},y_{k};z_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}H(x_{k+1},y_{k};z_{k})-\nabla_{y}H(x_{k},y_{k};z_{k}),y_{k+1}-y_{k}\rangle$
@@
where the first inequality follows from AM-GM inequality and the smoothness of $H$. Then, we can further get
@@
$+2\langle\nabla_{y}H(x(y_{k},z_{k}),y_{k};z_{k})-\nabla_{y}H(x_{k+1},y_{k};z_{k}),\;y_{k+1}-y_{k}\rangle$
@@
$\mathbb{E}\left[2\langle\nabla_{y}H(x(y_{k},z_{k}),y_{k};z_{k})-\nabla_{y}H(x_{k+1},y_{k};z_{k}),\;y_{k+1}-y_{k}\rangle\right]$ (28)
@@
where the first step follows from the smoothness of $H(x,y;z)$ and the third step follows from Lemma B.4. In the last step, we used the fact that $p>L$. Taking the expectation on both sides of (LABEL:eq:potential_bound1) and combining it with the bounds in (LABEL:eq:bound_term1) and (LABEL:eq:bound_term2), we can get
@@
Let $t_{1}=384pR_{y}\frac{1+\beta L+\beta L\sigma_{2}}{p-L}$. Then, using Lemma B.6, we have
@@
$\leq\;$ $384p\gamma\beta R_{y}\frac{1+\beta L+\beta L\sigma_{2}}{\beta(p-L)}\|y_{k}-y_{+}(z_{k})\|,$
@@
$\leq\;$ $384\gamma^{2}R_{y}\frac{1+\beta L+\beta L\sigma_{2}}{\beta(p-L)}\|y_{k}-y_{+}(z_{k})\|$
@@
$\leq\;$ $384p\alpha\gamma R_{y}\frac{1+\beta L+\beta L\sigma_{2}}{\beta(p-L)}\|y_{k}-y_{+}(z_{k})\|$
@@
By $\gamma\leq\frac{1}{\sqrt{KN}}$ and $\sum_{k=0}^{K-1}r_{k}^{2}\leq\frac{1}{b}$, we can get $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\mathfrak{g}(x_{k},y_{k})\|^{2}\right]\leq\mathcal{O}\left(\frac{N}{K}+\sqrt{\frac{N}{K}}\right)$, which leads to
@@
Substituting $\gamma\leq\frac{1}{\sqrt{KN}}$ and $\sum_{k=0}^{K-1}r_{k}^{2}\leq\frac{1}{b}$ into the above inequality, we can get
@@
Denote $\delta=\min_{j\in\mathcal{J}}((\overline{y})_{j}-(y)_{j})$ and $\overline{m}=\max_{j\in\mathcal{J}}(\overline{y})_{j}$. We also denote $\hat{y}=y+\beta c(x)$ and $\mathcal{P}_{j}[\cdot]=\mathcal{P}_{[0,(\overline{y})_{j}]}[\cdot]$ for simplicity. By the definitions of the KKT gap and stationarity measure, we have $\left\|\nabla_{x}f(x,y)\right\|=\|\mathfrak{g}_{x}(x,y)\|\leq\|\mathfrak{g}(x,y)\|.$
For the primal infeasibility and complementarity gap, we consider three cases for $(\hat{y})_{j}$: (1) $0\leq(\hat{y})_{j}\leq(\overline{y})_{j}$; (2) $(\hat{y})_{j}<0$; (3) $(\hat{y})_{j}>(\overline{y})_{j}$.
In the first case, we have $\mathcal{P}_{j}[(\hat{y})_{j}]=(\hat{y})_{j}$, so $(\mathfrak{g}_{y}(x,y))_{j}=\frac{1}{\beta}((y)_{j}-(\hat{y})_{j})=-c_{j}(x).$ Hence, we have $c_{j}(x)]_{+}\leq|c_{j}(x)|=|(\mathfrak{g}_{y}(x,y))_{j}|$ and $(y)_{j}|c_{j}(x)|\leq\overline{m}|(\mathfrak{g}_{y}(x,y))_{j}|$.
In the second case, we have $\mathcal{P}_{j}[(\hat{y})_{j}]=0$. Then, we can get $|(\mathfrak{g}_{y}(x,y))_{j}|=\frac{1}{\beta}(y)_{j}\geq 0$ and $(\hat{y})_{j}=(y)_{j}+\beta c_{j}(x)<0$, which leads to $c_{j}(x)<0$. Therefore, it yields $c_{j}(x)]_{+}=0$. Since $|c_{j}(x)|\leq\Lambda$ under Assumption 2.2, $(y)_{j}|c_{j}(x)|\leq\beta\Lambda|(\mathfrak{g}_{y}(x,y))_{j}|$.
In the third case, we have $\mathcal{P}_{j}[(\hat{y})_{j}]=(\overline{y})_{j}$ and $|(\mathfrak{g}_{y}(x,y))_{j}|=\frac{1}{\beta}((\overline{y})_{j}-(y)_{j})\geq\frac{\delta}{\beta}$. Since $(\hat{y})_{j}=(y)_{j}+\beta c_{j}(x)>(\overline{y})_{j}$, we can get $c_{j}(x)>\frac{\delta}{\beta}>0.$ Therefore, in this case $c_{j}(x)]_{+}=c_{j}(x)>0$. By Assumption 2.2, we have $c_{j}(x)]_{+}\leq\Lambda\leq\frac{\Lambda\beta}{\delta}|(\mathfrak{g}_{y}(x,y))_{j}|,$ and similarly $(y)_{j}|c_{j}(x)|\leq(\overline{y})_{j}\Lambda\leq\frac{\overline{m}\Lambda\beta}{\delta}|(\mathfrak{g}_{y}(x,y))_{j}|,$ where we used $|(\mathfrak{g}_{y}(x,y))_{j}|\geq\frac{\delta}{\beta}$.
Denote $\hat{c}=\max(1,\frac{\Lambda\beta}{\delta})$. Then, we have $c_{j}(x)]_{+}\leq\hat{c}|(\mathfrak{g}_{y}(x,y))_{j}|\leq\hat{c}\|\mathfrak{g}_{y}(x,y)\|$ and $(y)_{j}|c_{j}(x)|\leq\overline{m}\hat{c}|(\mathfrak{g}_{y}(x,y))_{j}|\leq\overline{m}\hat{c}\|\mathfrak{g}_{y}(x,y)\|$ Let $C=1+(\overline{m}+1)\hat{c}$. Finally, we can conclude $\mathcal{K}(x,y)\leq C\|\mathfrak{g}(x,y)\|$.